### v1

# thanks: Corresponding author

Abstract
In this paper, we systematically investigate the optical appearance of a Schwarzschild black hole illuminated by three geometrically thin accretion disk models under varying observational inclination angles. Based on the geometric relationship between the black hole and observer, we first divide the accretion disk into co-side and counter-side semi-disks. Subsequently, we analyze their differences in photon orbit distributions, transfer functions, and received intensities. Our results reveal distinct inclination-dependent behaviors between the two semi-disks: when the inclination approaches $\pi/2$, the co-side semi-disk develops richer ring structures with the lensed emission dominating the total brightness. Furthermore, through explicit specification of the emission profiles of the three models, we present optical images for both optically thin and thick disk scenarios at different inclinations. The results demonstrate that: (i) the bright rings in all three models become progressively compressed and deviate from circularity as inclination increases; (ii) for thick disks, partial rings are obscured and the overall intensity is lower than thin disks. These results may advance our understanding of general black hole imaging processes and provide a new approach to test gravitational theories through optical morphology studies.
I Introduction
As one of the most influential theories of gravity, general relativity (GR) has profoundly influenced our understanding of the nature of gravity. In weak-field regions, the validity of GR has been confirmed by numerous classical experiments [1]. Moreover, in strong-field regions, the successful detection of gravitational waves (GWs) from a binary black hole (BH) merger [2] and the release of the first image of the supermassive BH [3] provided further confirmation of GR’s predictions. The released BH image of M87* reveals a distinct bright ring encircling a central dark region. This luminous structure arises from photons that are deflected by the BH’s extreme gravity and redirected toward the observer, whereas the dark shadow area corresponds to photons that have fallen into the BH. The critical photon orbit (photon sphere) forms the boundary between the bright ring and the dark shadow. The shadow image of a BH reflects the properties of its surrounding spacetime geometry. By analyzing the shadow, one can extract key parameters of the BH [4, 5], which has drawn significant attention to BH shadow research. Synge was the first to theoretically analyze the formation of shadows for Schwarzschild BH [6]. Later, Bardeen investigated the shadow of Kerr BH [7], finding that the Kerr BH’s shadow deviates from the circular shape of Schwarzschild’s case, with the distortion increasing as the spin parameter grows. Since then, extensive theoretical studies have been conducted on BH shadows in various gravitational theories and under different matter-field couplings [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
Furthermore, considering astrophysical BHs in the universe, their surroundings inevitably contain various matter components. Under the influence of the BH’s strong gravitational field, this matter is accreted, forming different types of accretion structures around the BH [19]. Consequently, this leads to diverse BH shadow and bright ring configurations. Based on their geometric thickness and optical properties, accretion disks around BHs can be classified into thin and thick disks. The geometric thickness is typically quantified by the ratio of the disk thickness $D$ to its radius $R$ [20]. A disk is considered geometrically thin if $D/R\ll 1$. The optical thickness relates to the disk’s transparency. A disk is optically thick if it prevents the transmission of photons. Moreover, the distinct imaging features produced by different accretion disk models enable the verification of various gravitational theories through the optical appearance of BH accretion systems. In contrast to studies of geometrically thick disks [21], current research focus predominantly lies on geometrically thin disks. The first simulated image of a dynamically geometrically thin and optically thick accretion disk surrounding a Schwarzschild BH was produced by Luminet [22]. This pioneering work incorporated the observer’s inclination angle relative to the accretion disk, established angular relationships through spherical geometry, and ultimately generated a contour map on the observer’s plane displaying both the primary and secondary images of the BH. Subsequently, the study of BH imaging with such accretion disk models as emission sources, particularly under inclined viewing angles, has garnered widespread attention across various gravitational theories [20, 23, 24, 25, 26, 27, 28, 29, 30].
For optically thick accretion disks, the propagation of light is obstructed by the disk, which may lead to degradation of additional bright rings. Consequently, later studies adopted simpler geometrically thin and optically thin disk models. When investigating their imaging properties, Gralla et al. found that the observed image features are determined by the location and nature of emitting matter near the BH [31]. They further derived the optical appearance of three types of static, geometrically thin and optically thin accretion disks viewed face-on. Subsequent works have extensively investigated the luminous features of BHs illuminated by three static thin-disk models across various gravitational theories [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55].
In realistic astrophysical scenarios, observers rarely view BHs perfectly face-on to the equatorial plane, but substantial inclination angles exist. For example, the line of sight from Earth to M87* is not face-on its equatorial plane but maintains a specific angle [56, 3]. However, in current studies of these three static accretion disk models, systematic investigations of their BH images at varying inclination angles remain lacking. This motivates our focused investigation into the optical appearance of these three disk models under varying inclination conditions.
In this study, we examine three geometrically thin accretion disk models under both optically thin and optically thick conditions, and investigate the optical appearance characteristics of Schwarzschild BH illuminated by these distinct types of accretion disks. Specifically, we divide the accretion disk into co-side and counter-side semi-disks based on their positions relative to the observer and BH. Using this dual-semi-disk framework, we analyze the photon orbit number distributions and light-ray intersection patterns for both semi-disks at various inclinations. Then we introduce corresponding transfer functions to establish the positional relationship between the BH and observer’s plane, which are then used to calculate the received intensities. Finally, we obtain optical images for both thin and thick disk scenarios under the three static accretion disk models.
The structure of this paper is organized as follows. In Section II, we briefly recall photon trajectories around a Schwarzschild BH. In Section III, based on the relative positions between the BH and observer, we divide the accretion disk into two semi-disks for separate analysis, examining photon trajectory intersections with both semi-disks and introducing their respective transfer functions. In Section IV, we obtain the optical appearance of a Schwarzschild BH illuminated by three types of optically thin accretion disk models at various observational inclination angles. Section V presents the corresponding optical images for the optically thick disk scenarios of all three models, with systematic comparisons to their thin-disk counterparts. Finally, conclusions are provided in Section VI. Throughout this paper, we adopt geometric units with $G=c=1$ and set the BH mass $M=1$ for computational purposes.
II A Brief Recall of Photon Trajectories in Schwarzschild Black Hole Spacetime
The optical appearance of a Schwarzschild BH arises from photons near the BH being deflected and subsequently reaching the observer’s screen, forming an image. This implies that we need study the trajectories of photons around the BH to analyze its optical appearance. In Schwarzschild coordinates, the spacetime line element can be expressed as:
${\rm d}s^{2}=-f(r){\rm d}t^{2}+\frac{1}{f(r)}{\rm d}r^{2}+r^{2}\left({\rm d}\theta^{2}+\sin^{2}\theta{\rm d}\phi^{2}\right),$ (1)
where
$\begin{split}f(r)=1-\frac{2M}{r}.\end{split}$ (2)
Due to spherical symmetry, we restrict our analysis to photon trajectories confined to the equatorial plane ($\theta=\pi/2$). For photons, their geodesic tangent vectors $\left(\partial/\partial\tau\right)^{a}$ satisfy [57, 58]
$\begin{split}0:=g_{ab}\left(\frac{\partial}{\partial\tau}\right)^{a}\left(\frac{\partial}{\partial\tau}\right)^{b},\end{split}$ (3)
where $\tau$ is the affine parameter along the null geodesic. Within the Schwarzschild coordinate system, it follows that:
$\begin{split}0&=g_{ab}\left(\frac{\partial}{\partial\tau}\right)^{a}\left(\frac{\partial}{\partial\tau}\right)^{b}\\ &=-f(r)\left(\frac{{\rm d}t}{{\rm d}\tau}\right)^{2}+\frac{1}{f(r)}\left(\frac{{\rm d}r}{{\rm d}\tau}\right)^{2}+r^{2}\left(\frac{{\rm d}\phi}{{\rm d}\tau}\right)^{2}.\end{split}$ (4)
Note that in the Schwarzschild spacetime, there exist two Killing vector fields, $\left(\partial/\partial t\right)^{a}$ and $\left(\partial/\partial\phi\right)^{a}$, which correspond to the conserved quantities energy $E$ and angular momentum $L$, respectively. We have [57, 58]
$\begin{split}&E:=-g_{ab}\left(\frac{\partial}{\partial t}\right)^{a}\left(\frac{\partial}{\partial\tau}\right)^{b}=f(r)\frac{{\rm d}t}{{\rm d}\tau},\\ &L:=g_{ab}\left(\frac{\partial}{\partial\phi}\right)^{a}\left(\frac{\partial}{\partial\tau}\right)^{b}=r^{2}\frac{{\rm d}\phi}{{\rm d}\tau}.\end{split}$ (5)
Inserting Eq. (5) into Eq. (4) yields the photon trajectory equation, namely,
$\begin{split}\left(\frac{{\rm d}r}{{\rm d}\phi}\right)^{2}=r^{4}\left[\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}\right].\end{split}$ (6)
Here, $b\equiv L/E$ represents the photon’s impact parameter. For subsequent discussion, we reformulate the above expression using the substitution $u=1/r$, yielding
$\begin{split}\left(\frac{{\rm d}u}{{\rm d}\phi}\right)^{2}=\frac{1}{b^{2}}-f(\frac{1}{u})u^{2}.\end{split}$ (7)
Given a specific impact parameter $b$, its trajectory can be determined through Eq. (6). Figure  1 schematically displays the photon trajectories around the BH (represented by the black disk) for different impact parameters. Specifically, the red curve corresponds to photons with $b=b_{c}$ forming an unstable circular orbit around the BH due to gravitational deflection, where $b_{c}$ is referred to as the critical impact parameter. The black curves depict trajectories of photons with $b<b_{c}$ that ultimately fall into the event horizon, while the green curves represent photons with $b>b_{c}$ that are deflected by the BH and escape back to infinity.
Figure 1: The photon trajectories around a Schwarzschild BH. The black curves represent trajectories with impact parameters $b<b_{c}$, the red curve represents the photon trajectory corresponding to the critical impact parameter $b_{c}$, and the green curves depict trajectories with $b>b_{c}$. The black disk represents the BH.
Notably, due to spherical symmetry, photon trajectories around a Schwarzschild BH can always be considered as confined to a plane passing through the BH. In such a plane, the aforementioned unstable circular orbits exist. These unstable circular orbits collectively form an unstable photon sphere around the BH, where perturbed photons either fall into the BH or escape to infinity. Consequently, without considering emission from accretion disks, observers would detect a perfectly circular BH shadow silhouette from any viewing angle. However, when considering luminous accretion disks, the optical appearance of the BH becomes dependent on both the disk type and observer’s inclination angle. We now examine the optical appearance of BHs surrounded by accretion disks at various inclination angles.
III Trajectories of photons emitted from the thin accretion disk
III.1 Geometry of the black hole-observer system
For the Schwarzschild BH, we consider the presence of an optically thin and geometrically thin disk surrounding it. The disk emits radiation isotropically and is observed by distant observers at various inclination angles. The geometric relationship between the BH’s accretion disk and the observer is illustrated in Fig. 2.
Figure 2: The schematic diagram of the observer and BH coordinate systems, see [22, 29].
We assume the accretion disk lies strictly in the BH’s equatorial plane, with the observer positioned at infinity. The observer’s coordinate axis $O^{\prime}Y^{\prime}$ is coplanar with the BH’s axis $OM$, and the angle between $OZ$ and $OO^{\prime}$ defines the viewing inclination angle $\omega$. A photon emitted from point $P$ on the disk undergoes a deflection angle $\phi$ before reaching point $P^{\prime}$ on the observer’s plane. By projecting the observer’s plane $X^{\prime}O^{\prime}Y^{\prime}$ onto the BH coordinate system to obtain the $XOY^{\prime\prime}$ plane, we apply the spherical sine theorem to triangles $\triangle PXM^{\prime}$ and $\triangle PYM$ to establish the relationship between the deflection angle $\phi$, the viewing inclination $\omega$, and the polar angle $\eta$ in the observer’s coordinate system [22]:
$\cos{\phi}=\frac{\sin{\eta}}{\sqrt{\sin^{2}{\eta}+\cot^{2}{\omega}}}.$ (8)
Through further derivation, we obtain
$\phi=\arctan{(\frac{1}{\tan{\omega}\sin{\eta}})}.$ (9)
To facilitate subsequent discussion, we partition the BH’s accretion disk along the $XOY^{\prime\prime}$ plane into two distinct regions: the co-side semi-disk (located on the same side as the observer, to the right of the $XOY^{\prime\prime}$ plane) and the counter-side semi-disk (on the opposite side, to the left of the $XOY^{\prime\prime}$ plane).
III.2 Classification of photon trajectories
Photons emitted from the accretion disk and reaching the observer can equivalently be treated as backward-traced light rays from the observer intersecting with the disk near the BH. Each intersection corresponds to brightness contribution from the disk. Since light rays may intersect the accretion disk multiple times, we follow the conventional classification scheme of photon trajectories from [31], defining direct emission for rays intersecting the disk once, lensed emission for twice, and photon ring emission for three or more times by computing the orbit number $n=\phi/{(2\pi)}$, where $\phi$ represents the light deflection angle. Combining with Eq. (9), we derive:
$n=\frac{\phi}{2\pi}=\frac{1}{{2\pi}}\arctan{(\frac{1}{\tan{\omega}\sin{\eta}})}.$ (10)
Subsequently in Fig. 3, we demonstrate the intersection patterns between light rays (emitted from a distant observer at infinity on the right side) and the accretion disk around a Schwarzschild BH at non-zero inclination angles, specifically showing the counter-side semi-disk (left panel) and co-side semi-disk (right panel). The green dashed line $NY^{\prime\prime}$ represents the $XOY^{\prime\prime}$ plane from Fig. 2, while the blue solid line indicates the equatorial plane of the accretion disk. Black, orange, and red curves correspond to light rays intersecting the disk once, twice, and three or more times, respectively. We observe distinct intersection patterns between the counter-side semi-disk and co-side semi-disk, which motivates separate discussions of the orbit numbers $n_{1}$ for the counter-side semi-disk and $n_{2}$ for the co-side semi-disk.
Figure 3: Light trajectories from a distant right-side observer with inclination, intersecting the counter-side semi-disk (left) and co-side semi-disk (right) of a Schwarzschild BH. The blue solid curve represents the accretion disk, while the black, orange, and red curves correspond to light trajectories intersecting the disk once, twice, and three or more times, respectively.
In Figs. 4 and 5, we present the variations of orbit numbers $n_{1}$ and $n_{2}$ with impact parameter $b$ for different values of $\omega$ and $\eta$, respectively. In Fig. 4, with the polar angle fixed at $\eta=\pi/2$, the impact parameter range corresponding to the lensed emission region of the counter-side semi-disk gradually decreases with increasing viewing inclination, while that of the co-side semi-disk shows significant enlargement. Figure 5 demonstrates similar trends at fixed inclination $\omega=17\pi/36$ across various polar angles $\eta$. Additionally, we find the orbit numbers remain invariant under inclination changes when $\eta=0$.
We emphasize that for any given observer polar angle $\eta$, light rays intersect both semi-disks simultaneously as shown in Fig. 3. Moreover, the observed image exhibits exact symmetry about the $O^{\prime}Y^{\prime}$ axis. This symmetry reduces our analysis to the range $20\leq\eta\leq\pi/$ 2 while fully characterizing the entire image. All subsequent $\eta$ values are implicitly constrained to this angular domain without repeated specification.
Figure 4: The orbit numbers at fixed $\eta=\pi/2$ for varying inclination angles $\omega$. The first row shows the counter-side orbit numbers ($n_{1}$), while the second row displays the co-side orbit numbers ($n_{2}$). The black, orange, and red curves correspond to direct, lensed, and photon ring emissions respectively.
Figure 5: The orbit numbers at fixed $\omega=17\pi/36$ for varying inclination angles $\eta$. The first row shows the counter-side orbit numbers ($n_{1}$), while the second row displays the co-side orbit numbers ($n_{2}$). The black, orange, and red curves correspond to direct, lensed, and photon ring emissions respectively.
III.3 Transfer functions
To determine the optical appearance of a BH surrounded by a thin disk at various inclination angles, we must establish the mapping between each point on the accretion disk and its corresponding position on the observer’s plane. Here, we introduce the transfer function to characterize the relationship between points on the observer’s plane and their originating points on the accretion disk [31]. Notably, since light rays from the counter-side and co-side semi-disks undergo different deflection angles and cannot be treated uniformly, we classify and discuss the transfer functions separately.
For the the counter-side semi-disk, its transfer function $r_{\rm m}(b)$ can be expressed as:
$r_{\rm m}(b)=\frac{1}{u\left[\frac{m\pi}{2}+\arctan{({\tan{\omega}\sin{\eta}})},b\right]},\qquad m=1,2,3,\cdots,$ (11)
while for the co-side semi-disk, the transfer function $R_{\rm m}(b)$ takes the form:
$R_{\rm m}(b)=\frac{1}{u\left[\frac{m\pi}{2}-\arctan{({\tan{\omega}\sin{\eta}})},b\right]},\qquad m=1,2,3,\cdots.$ (12)
Here, $m$ denotes the radial position at the $m$-th intersection point between the light ray and the accretion disk. Notably, for the co-side transfer function, photons with $m=2$ actually intersect the counter-side semi-disk, yet their contributions appear in the co-side image. Therefore, the co-side (or counter-side) transfer functions essentially characterize the intersection positions on the accretion disk of light rays received at each point on the observer’s plane. Specifically, these transfer functions represent the contributions from different disk radii $r$ to the image. We derive these transfer functions by numerically solving Eq. 7.
Similar to previous analyses, Figs. 6 and 7 respectively present the transfer functions for the counter-side semi-disk and the first three transfer functions for the co-side semi-disk, with fixed $\eta=\pi/2$ and $\omega=17\pi/36$. Here, black, orange, and red curves correspond to $m=1$, $m=2$, and $m=3$ cases. The results align with prior findings on orbit numbers: as inclination $\omega$ (or polar angle $\eta$) increases, the impact parameter ranges for the second and third transfer functions of the counter-side semi-disk shrink, while those of the co-side semi-disk expand. Moreover, the slope of the first transfer function grows progressively with increasing $\omega$ and $\eta$, accompanied by a reduction in its impact parameter range–indicating contraction of the direct emission region on the observer’s plane at higher viewing angles. Notably, when the inclination approaches $\pi/2$, the lensed emission region of the co-side semi-disk becomes dozens of times larger than its counter-side counterpart. This implies that for moderately sized accretion disks, the co-side emission will be predominantly sourced from the lensed emission region.
Figure 6: Transfer functions at fixed $\eta=\pi/2$ for varying inclination angles $\omega$. The first row presents the counter-side transfer functions $r_{\rm m}(b)$, whereas the second row corresponds to the co-side transfer functions $R_{\rm m}(b)$. The black, orange, and red curves represent the transfer functions for $m=1$, $m=2$, and $m=3$ respectively.
Figure 7: Transfer functions at fixed $\omega=17\pi/36$ for varying inclination angles $\eta$. The first row presents the counter-side transfer functions $r_{\rm m}(b)$, whereas the second row corresponds to the co-side transfer functions $R_{\rm m}(b)$. The black, orange, and red curves represent the transfer functions for $m=1$, $m=2$, and $m=3$ respectively.
IV Optical Appearance of Black Hole at Different Inclinations with Thin Disks
In the previous section, we established the correspondence between emission points on the accretion disk and their projected positions on the observer’s plane. Incorporating both the local emission intensity of the optically thin disk and gravitational redshift effects, the received intensity at frequency $\nu_{o}$ originating from radius $r$ with emitted frequency $\nu_{e}$ is given by [31]:
$I_{\nu_{o}}^{\rm obs}=g^{3}I_{\nu_{e}}^{\rm em}(r),$ (13)
here $g$ denotes the redshift factor. The redshift factor can be calculated through $g=\frac{P_{\mu}u_{o}^{\mu}}{P_{\nu}u_{e}^{\nu}}$, where $P_{\mu}$ denotes the photon’s 4-momentum, while $u_{o}^{\mu}$ and $u_{e}^{\nu}$ correspond to the observer’s and the emitter’s 4-velocities [59], respectively. For a static observer at infinity, the 4-velocity is given by $u_{o}^{\mu}=(1,0,0,0)$, while for a static emitter, $u_{e}^{\nu}=(1/\sqrt{f(r)},0,0,0)$. The redshift factor is thus calculated as $g=\sqrt{f(r)}$. After integrating over the frequency $\nu_{e}$, we obtain:
[…]
The observed brightness originates from the intersections between light rays and the accretion disk. Therefore, we sum contributions from all intersection points on both the counter-side and co-side semi-disks to obtain the final received intensity:
$\begin{split}I_{\rm obs1}(b)&=\sum_{m}f(r)^{2}I_{\rm em}(r)|_{r=r_{\rm m}(b)},\\ I_{\rm obs2}(b)&=\sum_{m}f(r)^{2}I_{\rm em}(r)|_{r=R_{\rm m}(b)}.\end{split}$ (15)
Here, $I_{\rm obs1}(b)$ and $I_{\rm obs2}(b)$ represent the total received intensities from the counter-side and co-side semi-disks, respectively. As considered in [31], the $m=3$ transfer function has an extremely steep slope, contributing only to a drastically narrowed region of the image with negligible flux. We therefore exclude cases with $m>3$ in our analysis.
Thus, given a specific emission profile of the thin disk, we can derive the corresponding received intensity. We consider three commonly used emission models [31, 38]:
[…]
where $I_{0}$ is the maximum emission intensity. Additionally, $r_{\rm h}$, $r_{\rm ph}$, and $r_{\rm isco}$ denote the event horizon radius, photon sphere radius, and innermost stable circular orbit (ISCO) radius of the Schwarzschild BH, respectively. Their specific calculations are omitted here; see [46, 52] for details. Figure 8 presents the emission intensity profiles for three distinct models: Model-I, Model-II, and Model-III, which correspond to emission profiles initiating their decay at $r_{\rm isco}$, $r_{\rm ph}$, and $r_{\rm h}$, respectively.
Figure 8: The emission intensity profiles of the three models, where the red, blue, and green curves represent Model-I, Model-II, and Model-III, respectively.
Combining Eqs. (11), (12), and (15) we take Model-I as an example to demonstrate the received intensity on the observer’s plane at inclination angles of $\pi/4$ and $3617\pi/$ 36. Figures 9 and 10 show the intensity distributions for different polar angles ($\eta$) of the observer. From the figures, we observe that for the counter-side semi-disk, the number of intensity peaks in $I_{\rm obs1}$ shows no significant variation with polar angle $\eta$, while the distance between adjacent peaks displays only minor adjustments. Meanwhile, increasing the inclination angle reduces both the intensity at fixed $\eta$ and the peak separation distance. In contrast, the co-side intensity $I_{\rm obs2}$ exhibits more complex intensity profiles, developing additional emission peaks. For Model II and Model III, we briefly present the received intensity at select polar angles for inclination angles $\omega=\pi/4$ and $\omega=17\pi/36$ in Appendix A.
Figure 9: For Model-I at $\pi/4$ inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.
Figure 10: For Model-I at $3617\pi/$ 36 inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.
(a) Model-I:$\pi/36$
(b) Model-II:$\pi/36$
(c) Model-III:$\pi/36$
(d) Model-I:$18017\pi/$ 180
(e) Model-II:$18017\pi/$ 180
(f) Model-III:$18017\pi/$ 180
(g) Model-I:$\pi/4$
(h) Model-II:$\pi/4$
(i) Model-III:$\pi/4$
(j) Model-I:$17\pi/36$
(k) Model-II:$17\pi/36$
(l) Model-III:$17\pi/36$
Figure 11: Optical appearances of a Schwarzschild BH illuminated by three types of geometrically thin, optically thin accretion disks at varying inclination angles. Columns 1–3 display images for Model-I, Model-II, and Model-III at inclinations $\omega=\pi/36$, $18017\pi/$ 180, $\pi/4$, and $3617\pi/$ 36, respectively.
We then compute the received intensity for multiple polar angles at each inclination, and combine these results with the symmetry properties to project the intensity distribution onto a two-dimensional plane. This yields the final optical appearance of a thin-disk-illuminated Schwarzschild BH at various inclinations. As shown in Fig. 11, we present the optical appearances for three thin disk models at inclination angles of $\pi/36$, $18017\pi/$ 180, $\pi/4$, and $3617\pi/$ 36. And we find that as the inclination angle $\omega$ increases, the partial bright rings in the optical appearance of thin accretion disks for all three models become progressively compressed and deviate from circularity.
V Optical Appearance of Black Hole at Different Inclinations with Thick Disks
We now examine the optical appearance of a Schwarzschild BH surrounded by optically thick versions of the three accretion disk models. Unlike the optically thin case, the optically thick disk obstructs photon propagation, preventing some rays from reaching the observer and resulting in darker images. Therefore, we need to explicitly determine the spatial location of the thick accretion disk.
Specifically, we define the radial extent of the optically thick disk as:
  • Model-I: $r_{\rm isco}\leq r\leq 12$
  • Model-II: $r_{\rm ph}\leq r\leq 12$
  • Model-III: $r_{\rm h}\leq r\leq 12$
Additionally, we neglect self-irradiation effects (i.e., disk heating via internal photon absorption [22]), assuming time-independent emission profiles throughout.
For Model-I, the transfer functions of the counter-side semi-disk (similarly for the co-side) in Fig. 7 reveal that within the impact parameter range corresponding to $r_{\rm isco}\leq r_{1}(b)\leq 12$, the $m=1$ component dominates the image contribution. Due to disk obscuration, the $m=2$ and $m=3$ transfer functions contribute negligibly to the intensity in this region. Conversely, outside this impact parameter range where obscuration is absent, if $r_{2}(b)$ and $r_{3}(b)$ satisfy $r_{\rm isco}\leq r\leq 12$, the lensed ($m=2$) and photon ring ($m=3$) emission regions will contribute significantly to the image intensity. Thus, similar to the previous procedure, we fix the inclination angle $\omega$ and analyze the transfer functions for different observer polar angles $\eta$. We then calculate the received intensity using Eq. (15), considering only those transfer functions that actually contribute to the observed image.
The same computational procedure applies to Model-II and Model-III. However, since these models adopt disk radii starting from $r_{\rm ph}$ and $r_{\rm h}$ respectively, the lensed ($m=2$) and photon ring ($m=3$) contributions become negligible in regions where $r<r_{\rm ph}$ or $r<r_{\rm h}$. This allows for certain simplifications in our analysis.
Here, we take $\omega=17\pi/36$ as an example to demonstrate the optical appearance of a Schwarzschild BH illuminated by thick accretion disks under the three models, with direct comparison to the thin-disk cases in Fig. 12. The top row of Fig. 12 displays the images for the three thin-disk models, while the bottom row shows their thick-disk counterparts. Two key effects are evident: (i) the bright rings disappear in the thick-disk images due to disk obscuration, and (ii) the overall intensity of the thick disks is approximately $20\%$$40\%$ lower than that of the thin disks.
(a) Model-I:$17\pi/36$ (thin)
(b) Model-II:$17\pi/36$ (thin)
(c) Model-III:$17\pi/36$ (thin)
(d) Model-I:$17\pi/36$ (thick)
(e) Model-II:$17\pi/36$ (thick)
(f) Model-III:$17\pi/36$ (thick)
Figure 12: Optical appearances of the Schwarzschild BH illuminated by the three accretion disk models at an inclination angle of $\omega=17\pi/36$. The top row displays images for the three optically thin disk models, while the bottom row shows their optically thick counterparts.
VI Conclusion
In this paper, we investigate the optical appearance of a Schwarzschild BH surrounded by three geometrically thin and optically thin accretion disk models as observed from different inclination angles. By projecting the observer’s plane (the $XOY^{\prime\prime}$ plane in Fig. 2) onto the BH system, we divide the accretion disk into two hemispheres: the co-side semi-disk (facing the observer) and the counter-side semi-disk (opposite the observer), and this classification is consistently used throughout our analysis.
Following the methodology of [31], we categorize photon trajectories intersecting the disk once, twice, and three or more times as direct emission, lensed, and photon ring emissions, respectively. With this classification, each type of photon trajectory from both semi-disks produces distinct image components on the observer’s plane: the direct emission forms the primary image, while the lensed emission generates secondary images, and so forth. We then present the photon trajectories and orbit numbers for both semi-disks at fixed inclination $\omega$ and observer polar angle $\eta$. The results reveal striking differences between the two semi-disks: the lensed emission region is significantly larger on the co-side, and its size increases markedly with inclination, while the corresponding region on the counter-side shrinks under the same conditions.
To establish the relationship between the accretion disk and points on the observer’s plane, we introduce transfer functions $r_{\rm m}(b)$ and $R_{\rm m}(b)$ for the counter-side and co-side semi-disks, respectively. These functions characterize the disk intersection positions of light rays received at each point on the observer’s plane. The variation of these transfer functions with inclination angle follows the same trend as the orbit numbers. Notably, when both the inclination $\omega$ and polar angle $\eta$ approach $\pi/2$, the lensed emission region of the co-side semi-disk dominates the image contribution.
Subsequently, by specifying the emission profiles of the thin disk models and incorporating the transfer functions, we calculate the received intensity at different polar angles $\eta$ for a fixed inclination $\omega$. Our analysis reveals that the co-side semi-disk exhibits more complex peak structures in the received intensity across varying $\eta$, while the counter-side semi-disk maintains a nearly constant number of intensity peaks with only variations in their spacing. These results are then projected onto a two-dimensional plane to generate optical images for the three thin disk models at different inclinations, as shown in Fig. 11.
Finally, we extend our discussion to optically thick versions of the three models, comparing their images with the thin disk cases using $\omega=17\pi/36$ as a representative example. The results demonstrate that the thick disk’s obscuration effect eliminates certain bright rings and reduces the overall intensity by approximately $20\%$$40\%$ compared to thin disks.
In conclusion, this study extends the face-on imaging analysis of three thin accretion disk models to general inclination angles, better matching real astronomical observation conditions. We have obtained optical images for both thin and thick disk scenarios across various inclinations, enriching theoretical understanding and providing valuable references for future observational studies.
Acknowledgements.
This work is supported in part by NSFC Grant No. 12165005.
Appendix A Received intensities of Model-II and Model-III
In this appendix, we present the received intensities of Model-II and Model-III in Figs. 1316.
Figure 13: For Model-II at $\pi/4$ inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.
Figure 14: For Model-II at $3617\pi/$ 36 inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.
Figure 15: For Model-III at $\pi/4$ inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.
Figure 16: For Model-III at $3617\pi/$ 36 inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.